# NashOpt - A Python Library for Computing Generalized Nash Equilibria

### v1

@@ Abstract
NashOpt is an open-source Python library for computing and designing generalized Nash equilibria (GNEs) in noncooperative games with shared constraints and real-valued decision variables. The library exploits the joint Karush–Kuhn–Tucker (KKT) conditions of all players to handle both general nonlinear GNEs and linear–quadratic games, including their variational versions. Nonlinear games are solved via nonlinear least-squares formulations, relying on JAX for automatic differentiation. Linear–quadratic GNEs are reformulated as mixed-integer linear programs, enabling efficient computation of multiple equilibria. The framework also supports inverse-game and Stackelberg game-design problems. The capabilities of NashOpt are demonstrated through several examples, including noncooperative game-theoretic control problems of linear quadratic regulation and model predictive control. The library is available at https://github.com/bemporad/nashopt.
@@ 2 Problem Formulation
where $x_{i}\in\mathbb{R}^{n_{i}}$ collects the decision variables optimized by agent $i$, $i=1,\ldots,N$, $x_{-i}$ collects those of all other agents, $x=\mathop{\rm col}\nolimits(x_{1},\ldots,x_{N})\in\mathbb{R}^{n}$ is the stacked vector of all agents’ decision variables, $x^{\star}(p)$ is the generalized Nash equilibrium (GNE) of the game for a given parameter vector $p\in\mathbb{R}^{n_{p}}$, $n_{p}\geq 0$, where by $n_{p}=0$ we mean that no parameter enters (1). In the sequel, we will also write $x=(x_{i},x_{-i})$ to highlight the dependence of $x$ on $x_{i}$ and $x_{-i}$. In (1), $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is the objective of agent $i$, $g:\mathbb{R}^{n}\to\mathbb{R}^{n_{g}}$ and $h:\mathbb{R}^{n}\to\mathbb{R}^{n_{h}}$ define, respectively, (shared) inequality and equality constraints, and $\ell_{i}\in(\mathbb{R}\cup\{-\infty\})^{n_{i}}$, $u_{i}\in(\mathbb{R}\cup\{+\infty\})^{n_{i}}$ define the (local) box constraints for each player $i$. Moreover, $J:\mathbb{R}^{n}\times\mathbb{R}^{n_{p}}\to\mathbb{R}$ is the objective function used to shape the desired equilibrium behavior of the game, and $P\subseteq\mathbb{R}^{n_{p}}$ is a given feasible set for the parameter vector $p$, such as the hyper-box $P=\{p\in\mathbb{R}^{n_{p}}:\ \ell_{p}\leq p\leq u_{p}\}$ where $\ell_{p},u_{p}$ are (possibly infinite) bounds. Problem (1) is also referred to as an optimistic equilibrium of Single-Leader-Multi-Follower (Stackelberg) game [1], with the leader optimizing $p$ and the followers playing the GNEP parameterized by $p$.
A GNE $x^{\star}(p)$ is a vector where no agent can reduce their cost given the others’ strategies and feasibility constraints, i.e.,
$f_{i}(x_{i}^{\star},x_{-i}^{\star},p)\leq f_{i}(x_{i},x_{-i}^{\star},p)\qquad\forall x_{i}\in X_{i}(x_{-i}^{\star},p)$
where $X_{i}(x_{-i}^{\star},p)\subseteq\mathbb{R}^{n_{i}}$ is the feasible set of player $i$ defined by the constraints in (1) given the strategies $x_{-i}^{\star}(p)$ of the other players. We highlight special cases of (1) in the next sections.
@@ 2.1 Solving a specific GNEP
$\begin{array}[]{rl}x_{i}^{\star}=\arg\min_{\ell_{i}\leq x_{i}\leq u_{i}}&f_{i}(x_{i},p_{0})\\ \text{s.t. }&g(x,p_{0})\leq 0,\ h(x,p_{0})=0\\ &x_{-i}=x_{-i}^{\star},\quad i=1,\ldots,N.\end{array}$ (2)
@@ 2.3 Inverse game problem
Given an observed agents’ equilibrium $x_{\textrm{d}es}$ (or a desired one we wish to achieve), by setting $J(x^{\star},p)=\|x^{\star}-x_{\rm des}\|_{2}^{2}$, or $J(x^{\star},p)=\|x^{\star}-x_{\rm des}\|_{\infty}$ or $\|x^{\star}-x_{\rm des}\|_{1}$, we solve the inverse game of finding a vector $p$ (if one exists) such that $x^{\star}\approx x_{\textrm{d}es}$.
@@ 3 KKT Conditions and Nonlinear Least-Squares Formulation
Let us assume that $f_{i}$, $g$, and $h$ are continuously differentiable with respect to $x_{i}$ and that suitable constraint qualifications hold, such as the Linear Independence Constraint Qualification (LICQ) condition at the GNE $x^{\star}(p)$ [^1] The LICQ condition is verified if the gradients of the active constraints at $x^{\star}(p)$ $\bigl\{\nabla_{x_{i}}g_{j}(x^{\star},p):j\in\mathcal{A}(x^{\star})\bigr\}\;\cup\;\bigl\{\nabla_{x_{i}}h_{k}(x^{\star},p)\bigr\}\;\cup\;\bigl\{\pm e_{k}:x_{i,k}^{\star}=\ell_{i,k}\text{ or }u_{i,k}\bigr\}$ are linearly independent, where $e_{k}$ is the $k$-th unit vector of appropriate dimension, $\mathcal{A}(x^{\star})=\{j:g_{j}(x^{\star},p)=0\}$ is the set of active inequality constraints at $x^{\star}$, and $x_{i,k}$, $\ell_{i,k}$, and $u_{i,k}$ are the $k$-th components of $x_{i}$, $\ell_{i}$, and $u_{i}$, respectively.. Then, the necessary KKT conditions of optimality for each player $i$ read as follows [31, Theorem 12.1]:
@@
$\begin{array}[]{rcl}0&=&\nabla_{x_{i}}f_{i}(x,p)+\nabla_{x_{i}}g(x,p)^{\top}\lambda_{i}+\nabla_{x_{i}}h(x,p)^{\top}\mu_{i}\\ 0&=&h(x)\\ 0&=&\sqrt{\lambda_{i,j}^{2}+g_{j}(x,p)^{2}}-\lambda_{i,j}+g_{j}(x,p)\\ 0&=&\sqrt{v_{i,k}^{2}+(\ell_{i,k}-x_{i,k})^{2}}-v_{i,k}+\ell_{i,k}-x_{i,k}\\ 0&=&\sqrt{y_{i,k}^{2}+(x_{i,k}-u_{i,k})^{2}}-y_{i,k}+x_{i,k}-u_{i,k}\\ &&k=1,\ldots,n_{i},\ i=1,\dots,N,\ j=1,\ldots,n_{g}\end{array}$ (4)
where the last three residuals encode the primal feasibility (3c), dual feasibility (3d), and complementary slackness conditions (3e) using the Fischer–Burmeister nonlinear complementarity problem (NCP) function [17, 14] componentwise
$\phi(\alpha,\beta)=\sqrt{\alpha^{2}+\beta^{2}}-\alpha-\beta=0\quad\Longleftrightarrow\quad\alpha\beta=0,\ \alpha\geq 0,\ \beta\geq 0.$
@@
where $R(z,p)$ is the residual function defined in (4). In the sequel, we will denote by $\nu=\mathop{\rm col}\nolimits(\lambda,\mu,v,y)$ the stacked vector of all dual variables and $z=\mathop{\rm col}\nolimits(x,\nu)$ the overall vector of primal and dual variables. Given a solution $z^{\star}(p)$ satisfying (5), we will implicitly denote by $x^{\star}(p)$ the corresponding GNE is obtained by extracting from $z^{\star}(p)$ the components corresponding to the primal variables. We will also denote by $R(z)$ the residual function when $p$ is omitted or fixed.
For conditions of existence of a GNE for a given $p$, i.e., such as nonemptiness and compactness of the set of feasible decision vectors, convexity assumptions on $f$, $g$, and linearity of $h$, and constraint qualifications, we refer the reader to [16].
@@ 3.1 Determining a GNE via nonlinear least-squares
that serves as a merit function for the zero-finding problem (5[14, 11]. The nonlinear least-squares problem (6) can be solved via standard methods, such as Levenberg-Marquardt (LM) algorithms [27, 30], possibly exploiting sparsity in the Jacobian of $R(z,p_{0})$ with respect to $z$. If the minimum residual of the nonlinear least-squares problem (6) is zero (or, more realistically, below a given tolerance), then $z^{\star}(p_{0})$ contains a GNE solution $x^{\star}(p_{0})$ and the associated Lagrange multipliers for each agent’s problem associated with the parameter vector $p_{0}$.
@@ 3.3 Game design
where $z=\mathop{\rm col}\nolimits(x,\nu)$. Problem (7) can be solved via standard nonlinear programming methods, possibly exploiting sparsity in the Jacobian of $R(z,p)$ with respect to both $z$ and $p$.
@@ 3.4 Non-smooth regularization
When designing a parameter vector $p$, or finding an equilibrium among many existing ones for a given $p_{0}$, we may want to add the regularization term
@@
in the game-design objective $J$; for example, to find a sparse GNE solution $x^{\star}$ by choosing a sufficiently large $\alpha_{1}$. When $\alpha_{1}>0$, due to the non-differentiability of the $\ell_{1}$-norm at zero, we can reformulate the problem by splitting $x$ into positive and negative parts, $x=x_{p}-x_{m}$, with $x_{p},x_{m}\geq 0$ and apply the result in [5, Corollary 1] to minimize
$\begin{array}[]{rl}\min_{p,x_{p},x_{m},\nu}\quad&J(x_{p}-x_{m},p)+\frac{\rho}{2}\|R(\mathop{\rm col}\nolimits(x_{p}-x_{m},\nu),p)\|_{2}^{2}\\ &+\alpha_{1}\mathbf{1}^{\top}(x_{p}+x_{m})+\alpha_{2}(\|x_{p}\|_{2}^{2}+\|x_{m}\|_{2}^{2})\\ \text{s.t. }&p\in P\end{array}$ (9)
where $\nu$ collects all dual variables. In the special case we are just looking for a sparse GNE with no game-design objective (i.e., $J(x,p)\equiv 0$) and $\alpha_{2}>0$, by letting $\alpha_{3}\triangleq\sqrt{2\alpha_{2}}$, we have
@@
$\min_{p,x_{p},x_{m},\nu}$ $\frac{1}{2}\left\|\begin{array}[]{c}\alpha_{3}x_{p}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\\ \alpha_{3}x_{m}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\\ \sqrt{\rho}\ R(\mathop{\rm col}\nolimits(x_{p}-x_{m},\nu))\end{array}\right\|_{2}^{2}$ (10)
@@ 4 Linear Quadratic Games
Consider the following special case of (1) where the cost functions are quadratic, the constraints are linear, and the game design objective is convex piecewise affine:
@@
where $Q_{J}=Q_{J}^{\top}\succ 0$, $Q_{J}\in\mathbb{R}^{(n+n_{p})\times(n+n_{p})}$, $c_{J}\in\mathbb{R}^{n+n_{p}}$, or the sum of both, i.e.,
$J(x,p)=J_{\rm PWA}(x,p)+J_{\rm Q}(x,p)$ (12c)
@@
where $I_{\ell}$ and $I_{u}$ collect the rows of the identity matrix corresponding, respectively, to finite upper and lower bound, and $\ell_{\ell}$ and $u_{u}$ are the corresponding bounds.
The KKT conditions for each player’s optimization problem are necessary and sufficient for optimality due to the convexity of the cost functions and constraints and can be used to characterize the generalized Nash equilibria of the game:
@@
By stacking the KKT conditions (13) for all players, we can rewrite the generalized Nash equilibrium conditions as a single mixed-integer linear program (MILP) using standard big-M constraints to model the complementarity conditions (see, e.g., [24, Proposition 2]):
$\begin{array}[]{rl}\min_{p,x,\lambda,\mu,\delta,\sigma}\quad&\displaystyle{\sum_{j=1}^{n_{J}}\sigma_{j}+\frac{1}{2}[x^{T}\ p^{T}]Q_{J}\begin{bmatrix}x\\ p\end{bmatrix}+c_{J}^{T}\begin{bmatrix}x\\ p\end{bmatrix}}\\ \text{s.t. }&\sigma_{j}\geq D_{jk}x+E_{jk}p+h_{jk},\quad k=1,\dots,n_{j}\\ &Q^{i}_{i}x+c^{i}_{i}+F^{i}_{i}p+\bar{A}^{\top}_{i}\lambda_{i}+A_{\mathrm{eq,i}}^{\top}\mu_{i}=0\\ &\bar{A}x\leq\bar{b}+\bar{S}p\\ &A_{\mathrm{eq}}x=b_{\mathrm{eq}}+S_{\mathrm{eq}}p\\ &\lambda_{i}\geq 0,\ i=1,\ldots,N\\ &\bar{A}x-\bar{b}-\bar{S}p\leq M(1-\delta)\\ &\lambda_{i}\leq M\delta,\ i=1,\ldots,N\\ &\delta\in\{0,1\}^{m}\end{array}$ (14)
where $M\in\mathbb{R}$ is a sufficiently large constant, $\delta$ is a vector of binary variables introduced to model the complementarity conditions, and $\sigma$ is a vector of auxiliary variables used to define an upper-bound the objective function, $\sigma\in\mathbb{R}^{n_{J}}$. Clearly, at optimality $\sum_{i=1}^{n_{J}}\sigma_{j}=J_{\rm PWA}(x,p)$, as the first $n_{j}$ inequalities in (14) are all active. Note that using the same vector $\delta$ for all players ensures that shared constraints are active/inactive for all players simultaneously.
@@
4.1 Variational GNEs
As for the nonlinear case, variational GNEs can be obtained by enforcing that the Lagrange multipliers associated with the shared constraints are also the same for all players, i.e., by replacing $\{\lambda_{i}\}$ with a single vector $\lambda$ and $\{\mu_{i}\}$ with a single vector $\mu$, which further reduces the number of variables in (14).

### v2

@@ Abstract
NashOpt is an open-source Python library for computing and designing generalized Nash equilibria (GNEs) in noncooperative games with shared constraints and real-valued decision variables. The library exploits the joint Karush-Kuhn-Tucker (KKT) conditions of all players to handle both general nonlinear GNEs and linear–quadratic games, including their variational versions. Nonlinear games are solved via nonlinear least-squares formulations, relying on JAX for automatic differentiation. Linear–quadratic GNEs are reformulated as mixed-integer linear programs, enabling efficient computation of multiple equilibria. The framework also supports inverse-game and Stackelberg game-design problems. The capabilities of NashOpt are demonstrated through several examples, including noncooperative game-theoretic control problems of linear quadratic regulation and model predictive control. The library is available at https://github.com/bemporad/nashopt.
@@ 2 Problem Formulation
where $x_{i}\in\mathbb{R}^{n_{i}}$ collects the decision variables optimized by agent $i$, $i=1,\ldots,N$, $x_{-i}$ collects those of all other agents, $x=\mathop{\rm col}\nolimits(x_{1},\ldots,x_{N})\in\mathbb{R}^{n}$ is the stacked vector of all agents’ decision variables, $x^{\star}(p)$ is the generalized Nash equilibrium (GNE) of the game for a given parameter vector $p\in\mathbb{R}^{n_{p}}$, $n_{p}\geq 0$, where by $n_{p}=0$ we mean that no parameter enters (1). In (1), $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is the objective of agent $i$, $g:\mathbb{R}^{n}\to\mathbb{R}^{n_{g}}$ and $h:\mathbb{R}^{n}\to\mathbb{R}^{n_{h}}$ define, respectively, (shared) inequality and equality constraints, and $\ell_{i}\in(\mathbb{R}\cup\{-\infty\})^{n_{i}}$, $u_{i}\in(\mathbb{R}\cup\{+\infty\})^{n_{i}}$ define the (local) box constraints for each player $i$. Moreover, $J:\mathbb{R}^{n}\times\mathbb{R}^{n_{p}}\to\mathbb{R}$ is the objective function used to shape the desired equilibrium behavior of the game, and $P\subseteq\mathbb{R}^{n_{p}}$ is a given feasible set for the parameter vector $p$, such as the hyper-box $P=\{p\in\mathbb{R}^{n_{p}}:\ \ell_{p}\leq p\leq u_{p}\}$ where $\ell_{p},u_{p}$ are (possibly infinite) bounds. Problem (1) is also referred to as an optimistic equilibrium of Single-Leader-Multi-Follower (Stackelberg) game [1], with the leader optimizing $p$ and the followers playing the GNEP parameterized by $p$.
A GNE $x^{\star}(p)$ is a vector for which no agent can reduce their cost given the strategies of the other players and feasibility constraints, i.e.,
$f_{i}(\mathop{\rm col}\nolimits(x_{i}^{\star},x_{-i}^{\star}),p)\leq f_{i}(\mathop{\rm col}\nolimits(x_{i},x_{-i}^{\star}),p)\qquad\forall x_{i}\in X_{i}(x_{-i}^{\star},p)$
where, with a slight abuse of notation, we denoted by $\mathop{\rm col}\nolimits(x_{i},x_{-i})$ the vector obtained by collecting the components of $x_{i}$ and $x_{-i}$ in the right order, and $X_{i}(x_{-i}^{\star},p)\subseteq\mathbb{R}^{n_{i}}$ is the feasible set of player $i$ defined by the constraints in (1) given the strategies $x_{-i}^{\star}(p)$ of the other players. We highlight special cases of (1) in the next sections.
@@ 2.1 Solving a specific GNEP
$\begin{array}[]{rl}x_{i}^{\star}=\arg\min_{\ell_{i}\leq x_{i}\leq u_{i}}&f_{i}(x,p_{0})\\ \text{s.t. }&g(x,p_{0})\leq 0,\ h(x,p_{0})=0\\ &x_{-i}=x_{-i}^{\star},\quad i=1,\ldots,N.\end{array}$ (2)
@@ 2.3 Inverse game problem
Given an observed agents’ equilibrium $x_{\textrm{d}es}$ (or a desired one we wish to achieve), by setting $J(x^{\star},p)=\|x^{\star}-x_{\rm des}\|_{2}^{2}$ (or $J(x^{\star},p)=\|x^{\star}-x_{\rm des}\|_{\infty}$, $\|x^{\star}-x_{\rm des}\|_{1}$), we solve the inverse game of finding a vector $p$ (if one exists) such that $x^{\star}\approx x_{\textrm{d}es}$.
@@ 3 KKT Conditions and Nonlinear Least-Squares Formulation
Let us assume that $f_{i}$, $g$, and $h$ are continuously differentiable with respect to $x_{i}$ and that suitable constraint qualifications hold, such as the Linear Independence Constraint Qualification (LICQ) condition at the GNE $x^{\star}(p)$ [^1] The LICQ condition is verified if the gradients of the active constraints at $x^{\star}(p)$ $\bigl\{\nabla_{x_{i}}g_{j}(x^{\star},p):j\in\mathcal{A}(x^{\star})\bigr\}\;\cup\;\bigl\{\nabla_{x_{i}}h_{k}(x^{\star},p)\bigr\}\;\cup\;\bigl\{\pm e_{k}:x_{i,k}^{\star}=\ell_{i,k}\text{ or }u_{i,k}\bigr\}$ are linearly independent, where $e_{k}$ is the $k$-th unit vector of appropriate dimension, $\mathcal{A}(x^{\star})=\{j:g_{j}(x^{\star},p)=0\}$ is the set of active inequality constraints at $x^{\star}$, and $x_{i,k}$, $\ell_{i,k}$, and $u_{i,k}$ are the $k$-th components of $x_{i}$, $\ell_{i}$, and $u_{i}$, respectively.. Then, the necessary KKT conditions of optimality for each player $i$ read as follows [32, Theorem 12.1]:
@@
$\begin{array}[]{rcl}0&=&\nabla_{x_{i}}f_{i}(x,p)+\nabla_{x_{i}}g(x,p)^{\top}\lambda_{i}+\nabla_{x_{i}}h(x,p)^{\top}\mu_{i}\\ 0&=&h(x,p)\\ 0&=&\sqrt{\lambda_{i,j}^{2}+g_{j}(x,p)^{2}}-\lambda_{i,j}+g_{j}(x,p)\\ 0&=&\sqrt{v_{i,k}^{2}+(\ell_{i,k}-x_{i,k})^{2}}-v_{i,k}+\ell_{i,k}-x_{i,k}\\ 0&=&\sqrt{y_{i,k}^{2}+(x_{i,k}-u_{i,k})^{2}}-y_{i,k}+x_{i,k}-u_{i,k}\\ &&k=1,\ldots,n_{i},\ i=1,\dots,N,\ j=1,\ldots,n_{g}\end{array}$ (4)
where the last three residuals encode the primal feasibility (3c), dual feasibility (3d), and complementary slackness conditions (3e) using the Fischer–Burmeister nonlinear complementarity problem (NCP) function $\phi:\mathbb{R}\to\mathbb{R}$ [18, 15] componentwise
$\phi(\alpha,\beta)\triangleq\sqrt{\alpha^{2}+\beta^{2}}-\alpha-\beta=0\quad\Longleftrightarrow\quad\alpha\beta=0,\ \alpha\geq 0,\ \beta\geq 0.$
@@
where $R(z,p)$ is the residual function defined in (4). In the sequel, we will denote by $\nu=\mathop{\rm col}\nolimits(\lambda,\mu,v,y)$ the stacked vector of all dual variables, so that $z=\mathop{\rm col}\nolimits(x,\nu)$ is the overall vector of primal and dual variables. Given a solution $z^{\star}(p)$ satisfying (5), we will implicitly denote by $x^{\star}(p)$ the corresponding GNE is obtained by extracting from $z^{\star}(p)$ the components corresponding to the primal variables. We will also denote by $R(z)$ the residual function when $p$ is omitted or fixed.
For conditions of existence of a GNE for a given $p$, such as nonemptiness and compactness of the set of feasible decision vectors, convexity assumptions on $f$, $g$, and linearity of $h$, and constraint qualifications, we refer the reader to [17].
@@ 3.1 Determining a GNE via nonlinear least-squares
that serves as a merit function for the zero-finding problem (5[15, 12]. The nonlinear least-squares problem (6) can be solved via standard methods, such as Levenberg-Marquardt (LM) algorithms [28, 31], possibly exploiting sparsity in the Jacobian of $R(z,p_{0})$ with respect to $z$. If the minimum residual $R(z^{\star}(p_{0}),p_{0})$ of the nonlinear least-squares problem (6) is zero (or, more realistically, below a given positive tolerance), then $z^{\star}(p_{0})$ contains a GNE solution $x^{\star}(p_{0})$ and the associated Lagrange multipliers for each agent’s problem associated with the parameter vector $p_{0}$.
@@ 3.3 Game design
where $z=\mathop{\rm col}\nolimits(x,\nu)$. Problem (7) can be solved via standard nonlinear programming methods with respect to both $z$ and $p$.
@@ 3.4 Non-smooth regularization
When designing a parameter vector $p$, or finding an equilibrium among many existing ones for a given $p_{0}$, we can add the regularization term
@@
in the game-design objective $J$; for example, to find a sparse GNE solution $x^{\star}$, by choosing a sufficiently large $\alpha_{1}$. When $\alpha_{1}>0$, due to the non-differentiability of the $\ell_{1}$-norm at zero, we can reformulate the problem by splitting $x$ into positive and negative parts, $x=x_{p}-x_{m}$, with $x_{p},x_{m}\geq 0$ and apply the result in [5, Corollary 1] to minimize
$\begin{array}[]{rl}\min_{p,x_{p},x_{m},\nu}\quad&J(x_{p}-x_{m},p)+\frac{\rho}{2}\|R(\mathop{\rm col}\nolimits(x_{p}-x_{m},\nu),p)\|_{2}^{2}\\ &+\alpha_{1}\mathbf{1}^{\top}(x_{p}+x_{m})+\alpha_{2}(\|x_{p}\|_{2}^{2}+\|x_{m}\|_{2}^{2})\\ \text{s.t. }&p\in P,\quad x_{p}\geq 0,\ x_{m}\geq 0\end{array}$ (9)
where $\nu$ collects all dual variables.
In the special case we are just looking for a sparse GNE with no game-design objective (i.e., $J(x,p)\equiv 0$) and $\alpha_{2}>0$, by letting $\alpha_{3}\triangleq\sqrt{2\alpha_{2}}$, we have
@@
$\min_{p,x_{p},x_{m},\nu}$ $\frac{1}{2}\left\|\begin{bmatrix}\alpha_{3}x_{p}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\\ \alpha_{3}x_{m}+\frac{\alpha_{1}}{\alpha_{3}}\mathbf{1}\\ \sqrt{\rho}\ R(\mathop{\rm col}\nolimits(x_{p}-x_{m},\nu))\end{bmatrix}\right\|_{2}^{2}$ (10)
@@
The approach can be immediately extended to the case $J(x,p)=\frac{1}{2}\|f_{J}(x,p)\|_{2}^{2}$ for some vector function $f_{J}:\mathbb{R}^{n}\times\mathbb{R}^{n_{p}}\to n_{f}$ by extending the residual vector in (10) with $f_{J}(x_{p}-x_{m},p)$.
@@ 4 Linear Quadratic Games
Consider the following special case of (1) where the agents’ cost functions are quadratic and the constraints are linear:
@@
where $Q_{J}=Q_{J}^{\top}\succeq 0$, $Q_{J}\neq 0$, $Q_{J}\in\mathbb{R}^{(n+n_{p})\times(n+n_{p})}$, $c_{J}\in\mathbb{R}^{n+n_{p}}$, or the sum of both, i.e.,
$J(x,p)=J_{\rm PWA}(x,p)+J_{\rm Q}(x,p).$ (12c)
@@
where $I_{\ell}$ and $I_{u}$ collect the rows of the identity matrix corresponding, respectively, to finite upper and lower bound, and $\ell_{\ell}$ and $u_{u}$ are the corresponding bounds. Let $m$ be the total number of resulting inequality constraints.
The KKT conditions for each player’s optimization problem are necessary and sufficient for optimality, due to the convexity of the cost functions and constraints, and can be used to characterize the generalized Nash equilibria of the game:
@@
By stacking the KKT conditions (13) for all players, we can rewrite the generalized Nash equilibrium conditions as a single mixed-integer linear program (MILP) using standard big-M constraints to model the complementarity conditions (see, e.g., [25, Proposition 2]):
$\begin{array}[]{rl}\min_{p,x,\lambda,\mu,\delta,\sigma}\quad&\displaystyle{\frac{1}{2}[x^{T}\ p^{T}]Q_{J}\begin{bmatrix}x\\ p\end{bmatrix}+c_{J}^{T}\begin{bmatrix}x\\ p\end{bmatrix}}+\sum_{j=1}^{n_{J}}\sigma_{j}\\ \text{s.t. }&\sigma_{j}\geq D_{jk}x+E_{jk}p+h_{jk},\quad k=1,\dots,n_{j}\\ &Q^{i}_{i}x+c^{i}_{i}+F^{i}_{i}p+\bar{A}^{\top}_{i}\lambda_{i}+A_{\mathrm{eq},i}^{\top}\mu_{i}=0\\ &\bar{A}x\leq\bar{b}+\bar{S}p\\ &A_{\mathrm{eq}}x=b_{\mathrm{eq}}+S_{\mathrm{eq}}p\\ &\lambda_{i}\geq 0,\ i=1,\ldots,N\\ &\bar{A}x-\bar{b}-\bar{S}p\leq M(1-\delta)\\ &\lambda_{i}\leq M\delta,\ i=1,\ldots,N\\ &\delta\in\{0,1\}^{m}\end{array}$ (14)
where $M\in\mathbb{R}$ is a sufficiently large positive constant, $\delta$ is a vector of binary variables introduced to model the complementarity conditions, and $\sigma$ is a vector of auxiliary variables used to define an upper-bound the objective function, $\sigma\in\mathbb{R}^{n_{J}}$. Clearly, at optimality $\sum_{i=1}^{n_{J}}\sigma_{j}=J_{\rm PWA}(x,p)$, as the first $n_{j}$ inequalities in (14) become all active. Note that using the same vector $\delta$ for all players ensures that shared constraints are active/inactive for all players simultaneously.